OpenStudy (anonymous):

A RECTANGULAR parking lot is 5over6 mile long, and 3over4 mile wide, how much fencing is needed to enclose the lot? use the formula P= 2L+2W..

OpenStudy (anonymous):

By 5over6 do you mean 5/6 and same for 3over4?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

Okay, so basically, you're just gonna plug everything into your equation. P=2L+2W P=2(5/6)+2(3/4) Can you solve from there or should I keep going?

OpenStudy (anonymous):

keep goin

OpenStudy (anonymous):

Okay. P=2L+2W P=2(5/6)+2(3/4) P=(5/3)+(3/2) P=19/6

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

You're very welcome